Calculate the activation parameters; Ea, H‡
, and S‡
at 100
o C from the data given for the
reaction below.
Temperature oC
k x 104
(s-1)
60.0
0.30
70.0
0.97
75.0
1.79
80.0
3.09
90.0
8.92
95.0
15.90
1
Expert's answer
2020-10-01T07:54:46-0400

Using the Arrhenius equation, the activation energy "E_a" can be calculated from the plot of "\text{ln}(k)" in function of "1\/T" :

"\text{ln}(k) = \text{ln}(A) - \frac{E_a}{RT}" ,

where "A" is the frequency factor constant and "R" is the gas constant, 8.314 J/(mol K).


The activation energy is calculated from the slope of the trend line:

"E_a = 13862\cdot8.314 = 115.25" kJ/mol.

Now we can get the standard enthalpy of activation "\Delta H^‡" at 100°C, using the following relation:

"\Delta H^‡ = E_a-RT"

"\Delta H^‡ = 115249-8.314\cdot(273.15+100) = 112.15" kJ/mol.

However, in order to get the standard entropy of activation, we must use Eyring equation:

"\text{ln}(\frac{k}{T}) = \text{ln}(\frac{k_B}{h}) +\frac{\Delta S^‡}{R} - \frac{\Delta H^‡}{RT}" .

Using its linearization, the slope will be "-\Delta H^‡/R" , while the intercept will be a fuction of the entropy: "\Delta S^‡/R + \text{ln}(\frac{k_B}{h})" , where "k_B" is the Boltzmann constant 1.381 10-23 J/K, and "h" is the Plank constant, 6.626 10-34 Js.



From these relations, one can get both the enthalpy "\Delta H^‡ = 13511\cdot8.314 = 112.33" kJ/mol and the entropy of activation "\Delta S^‡ = (24.317-23.76)\cdot8.314 = 4.60" (J/(mol K)). Note that the enthalpies obtained are very close: 112.33 and 112.15 kJ/mol.

Answer: "E_a = 115.25" kJ/mol, "\Delta H^‡ =112.33" (112.15) kJ/mol and "\Delta S^‡ = 4.60" J/(mol K).